OpenStudy (anonymous):

Verify that x-1 is a factor of x^(n) - 1

hartnn (hartnn):

if x-a is the factor of f(x) then f(a) = 0 now put x=1 in x^n-1 and see whether u get answer as 0.

OpenStudy (anonymous):

put x=1 in above equation and if it reduces to zero then thats the factor

OpenStudy (anonymous):

x^n - 1 = (x-1)(1 + x + x^2 +...+ x^n-1) Would you say that this is a good verification?

OpenStudy (anonymous):

It just seemed too easy to me previously lol

hartnn (hartnn):

yup, that verification will also do. just put there x=1 so (1-1)(....) = 0(...) = 0

OpenStudy (anonymous):

Got it! Thank you!

hartnn (hartnn):

welcome :)